OpenStudy (anonymous):

Find the vertical asymptotes, if any, of the graph of the rational function. f(x) = x-4/x(x-4)

OpenStudy (anonymous):



OpenStudy (anonymous):

Vertical asymptotes means that the denominator is equal to zero.

OpenStudy (anonymous):

you factor the denominator right

OpenStudy (anonymous):

yes but see, x-4, you would say that x=4 is a pol, but it's not, because you can cancel it with the numerator.

OpenStudy (anonymous):

\[\frac{(x-4)}{x(x-4)}=\frac{1}{x} \]